OpenStudy (anonymous):

Is 2^n + 3^n (where n is an integer) ever the square of a rational number?

OpenStudy (anonymous):

2^n + 3^n=x^2

OpenStudy (anonymous):

So yes or no?

OpenStudy (anonymous):

If you agree with my equation than yes :P

OpenStudy (anonymous):

Timo, it is a bit more complicated than that. The theorem as posed is: \[\exists n\in\mathbb{N} \text{ such that }\sqrt{2^n+3^n}\in\mathbb{Q}.\]As for the answer, I don't know.

OpenStudy (dumbcow):

it is highly unlikely...but im not sure how to mathematically prove it either way

OpenStudy (anonymous):

woah that looks confusing:/ It was a question asked by my friend but i don't get any of this really.

OpenStudy (dumbcow):

i could run a program...but theoretically there are an infinite possibilities for n so you couldn't say never...maybe proof by induction

OpenStudy (anonymous):

I thought induction at first too, by contradiction, but I think that's a blind alley. @jazy, it's a great question, and is fun to think about. I don't have the tools yet to answer it well, other than to write a program like dumbcow suggested. I'm sure there are ways of thinking about the question that I don't have access to yet that would make it answerable.

OpenStudy (anonymous):

I looked it up on google and came accross this: http://www.qbyte.org/puzzles/p155s.html I guess thats where she got it from....always trying to trick me!

OpenStudy (anonymous):

Yep, I was just looking it up on google as well. Turns out the proof uses modular arithmetic, which I would not have thought of. Cool stuff.

OpenStudy (anonymous):

Yup, I guess it's pretty cool to think about...Once I get there:)